find the value for z and w for 4z-1 = -2iw


1
Expert's answer
2021-08-19T08:24:21-0400
"4z-1=-2iw"

"4z=1-2iw"

"z=\dfrac{1}{4}-\dfrac{1}{2}iw"

"w=-\dfrac{1}{2}i+2iz"